y varies directly as the square of x. When y is 36, x is 2. Find y when x is 3.
Answer by venugopalramana(3286)

y varies directly as the square of x. When y is 36, x is 2. Find y when x is 3.
Y VARIES AS X OR VARIES DIRECTLY AS X MEANS Y=K*X WHERE K IS A CONSTANT....
Y VARIES INVERSELY AS X MEANS Y=K*(1/X)=K/X...WHERE K IS A CONSTANT..USEING THIS WE GET FOR DIRECT VARIATION
Y=KX
WHEN X=2,Y=36 SUBSTITUTING TO GET K ,WE HAVE
36 =2K...OR K=36/2=18
SO SINCE K IS A CONSTANT WE HAVE Y=18X..SO WHEN X=3,WE HAVE
Y=18X=18*3=54